# Inferring hyperuniformity from local structures via persistent homology

### v1

@@ 7 Discussion and Conclusions
We investigated whether the HU character of a point pattern can be deduced from local structures as described by the persistent homology and how such a task can be performed. The results show that the information contained in persistence diagrams allows for detecting the HU character as well as extracting and exploiting further details.
@@
Quantitative comparisons between persistence diagrams are also shown to enable the estimation of global properties, as shown in terms of the parameters entering the (ideal) structure factor of the point pattern from which persistence diagrams are computed. We remark that the HU character can be, in general, easily computed from equation (3) by looking at the power law decay for small $\mathbf{k}$ as well as $\widetilde{H}$, in infinite or periodic systems. For non-periodic finite size systems, either extrapolation or improved statistical estimations [39] should be considered, although defect perturbations and finite size effects induce additional phenomenology (see, e.g., [40, 41, 42, 43]), which can be hardly fully captured by global measures only. In these cases, estimating effective structure factor parametrization based on local structures proposed here may thus represent a practical method to characterize the system comprehensively, accounting for additional effects from small- to intermediate-scales. In addition, applications of ML methods and reference persistence diagrams can be envisaged to screen large datasets efficiently.
@@
Given that we show that HU arrangements entail peculiar local arrangements and/or geometrical properties, we envision a detailed inspection of the dependence of physical properties on the latter, for instance, in the transient phase of transport phenomena in porous media [44] or diffusion [45]. It will be essential to generate a pattern with targeted topological features independently from global constraints to study how related effects can be decoupled – augmenting the design space – and eventually fully characterizing their effects. It is in this context that the method we proposed to generate (point) patterns given the topological features encoded in persistence diagrams is particularly interesting, besides addressing a theoretical, fundamental aspect. We remark that the developed method may find application well beyond HU patterns and in the more general context of persistent homology.
@@
A natural direction for future research consists of extending the proposed analysis and the study of persistent homology in general to scalar and vectorial fields exhibiting hyperuniformity. This also includes inspecting different parametrizations of the structure factor to study specific physical systems. Moreover, the aforementioned physical properties can be targeted by devising inverse methods for the generation of HU arrangements entailing specific properties. In this regard, we have here shown that features based on persistent homology well correspond to HU characters. With this work, we have then provided crucial information for their exploitation as descriptors for inverse and data-driven inverse design concepts.
@@ Appendix B Standard deviation of the accuracy rates by ML models
Appendix C Details of the generation of point pattern from persistence diagrams

### v2

@@ 8 Discussion and Conclusions
We investigated whether local arrangements realized in point patterns can be used to deduce HU or non-HU characters. It is shown that inverse relations between hyperuniformity and persistent homology can be inverted in an approximated fashion. This is achieved by leveraging persistent homology and data-driven techniques, and we illustrate how such a task can be performed. The results show that the information contained in persistence diagrams allows for characterizing and using local (geometrical and topological) features and, in general, can be used to detect arrangements analogous to the ones realized in HU patterns. We envision this latter aspect to be crucial for characterizing systems with the emergence of hyperuniformity in confined regions.
@@
Quantitative comparisons between persistence diagrams allow for the estimation of global properties, as shown in terms of the parameters entering the (ideal) structure factor of the point pattern from which persistence diagrams are computed. We remark that the HU character can be, in general, easily computed from equation (3) by looking at the power law decay for small $\mathbf{k}$ as well as $\widetilde{H}$, in infinite or periodic systems. In this regard, the proposed framework can be used as a detection method, but the direct calculation of the structure factor (or the number variance) would be more accurate and efficient. The estimation of the parameters discussed here, however, shows that information contained in distributions of local structures can be used for quantitative estimations.
@@
More interesting is the detection of traits of hyperuniformity in non-periodic, finite-size systems, e.g., where the underlying physics may enforce a HU character (e.g., the physical systems in Ref. [10, 7, 21]) while finite domain size may disrupt the ideal HU behavior. The latter is indeed rigorously defined for infinitely extended domains only (and mimicked in periodic domains). For finite-size systems, either extrapolation or improved statistical estimations [50] should be considered, although defect perturbations and finite size effects induce additional phenomenology (see, e.g., [51, 52, 53, 54]), which can be hardly fully captured by global measures only. In these cases, estimating effective structure factor parametrization based on local structures proposed here may thus represent a practical method to characterize the system comprehensively, accounting for additional effects from small- to intermediate scales. In addition, applications of ML methods and reference persistence diagrams can be envisaged to screen large datasets efficiently.
We recall that our analysis builds on the model structure factor introduced in Eq. (5). Besides introducing a specific, tunable HU (or non-HU) character at large length scales, it enforces the same randomness for all the patterns at small scales. We expect that the considered framework based on data analysis and advanced statistical methods, namely persistent homology and ML methods, would benefit from any additional information for small scales ($|\mathbf{k}|\,$$>$$\,K$), such as other functional forms or even an unprescribed structure factor –thus adapting in an uncontrolled manner to global order– for the corresponding range of wave numbers. As the analysis of the distribution of local structures proposed here is done by constraining the pattern to feature ideal randomness at short scales, we expect that the result reported here represents a lower bound in terms of accuracy for other strongly correlated HU systems.
@@
Given that we show that HU arrangements entail peculiar local arrangements and/or geometrical properties, we envision a detailed inspection of the dependence of physical properties on the latter, for instance, in the transient phase of transport phenomena in porous media [55] or diffusion [30]. It will be essential to generate a pattern with targeted topological features independently from global constraints to study how related effects can be decoupled, thus augmenting the design space. It is in this context that the method we proposed to generate (point) patterns given the topological features encoded in persistence diagrams is particularly interesting, besides addressing a theoretical, fundamental aspect. We remark that the developed method may find application well beyond HU patterns and in the more general context of persistent homology.
@@
A natural direction for future research consists of extending the proposed analysis and the study of persistent homology to scalar and vectorial fields exhibiting hyperuniformity. Examples are two-phase media obtained by decorating point patterns [29] with circles or other shapes as well as considering parametrization for prototypical HU fields such as spinodal patterns [19]. This also includes inspecting different parametrizations of the structure factor to study specific physical systems. Inverse design methods for generating HU arrangements entailing specific properties will also be explored. In this regard, we have shown that features based on persistent homology correspond well to HU characters. With this work, we have then provided crucial information for their exploitation as descriptors for inverse and data-driven inverse design concepts.
@@ Appendix B Standard deviation of the accuracy rates by ML models
Appendix C Estimation of $H$ with Neural Network regression
While we focused in section 4 on classification between hyperuniform and non-hyperuniform patterns given a threshold on the parameter $H$, this section shows that the actual value of parameter $H$ can be estimated from the persistence diagram addressing a regression problem.
It is well known in ML studies that regression generally necessitates larger training data than classification. For the following, we consequently simplify our problem by considering solely patterns with fixed $\alpha$ and $K$, but with a random $H$ value between $5\cdot 10^{-4}$ and $1$ (i.e., not restricted to the values in (7)). Using the feature derived from the Wasserstein distance presented in 4.1.3, a neural network with three hidden layers of respectively $128$, $128$ and $256$ neurons is trained over $584$ realizations of such patterns with $\alpha=10$ and $K=45\pi$. This neural network was implemented with the function MPLRegressor of the python library sklearn. After testing the model over $146$ realizations, the accuracy of the regression is effectively good, as summarized in figure 13. It is noteworthy that larger deviations are observed for a nominal $H$ smaller than $10^{-2}$: this remark is consistent with the fact that the local geometry of arrangements is only barely distinguishable for $H<10^{-2}$, as mentioned in [23] and recalled in section 2.
Figure 13: Estimation of the parameter $H$ via a neural network regression for fixed $\alpha=10$ and $K=45\pi$. Each cross represents one estimation. The accuracy is very good for nominal values higher than $10^{-2}$ while larger deviations are observed below this threshold, which is consistent with observations made in [23]. This threshold is indicated with dashed grey lines.
Appendix D Details of the generation of point pattern from persistence diagrams